OpenStudy (anonymous):

The equation of a parabola on an xy plane is xy = -6. By what angle must the coordinate axes be rotated to make the transverse axis horizontal? 30° 45° 60° 90° 0°

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@King.Void.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (amistre64):

thats not a parabola, its a hyperbola ... if memory serves

OpenStudy (amistre64):

y = 1/x is a hyperbola

OpenStudy (anonymous):

well it says a parabola

OpenStudy (anonymous):

It probably is a hyperbola the wording just wrong

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y+%3D+-6%2Fx it is clearly a hyperbola since it dos not even conform to the definition of a parabola

OpenStudy (amistre64):

transverse, which axis is that?

OpenStudy (anonymous):

yea I wolfram it too

OpenStudy (anonymous):

insn't on the x axis?

OpenStudy (amistre64):

transverse, the axis that the object crosses ... it should currently be along the y= -x line

OpenStudy (anonymous):

yea that too

OpenStudy (amistre64):



OpenStudy (anonymous):

so that is a 30 degree angle?

OpenStudy (anonymous):

Ima jsut go with 30 degrees since you leave

OpenStudy (amistre64):

i moderate the site, i have to leave at times ..

OpenStudy (amistre64):

y = 1/x and y= 6/x are the same 'axial' representations of the hyperbola what degree is y = -x from the x axis?